what is the slope of each line and then determine whether the two lines are parallel, perpendicular or neither
-5x+3y=6
-20x+12y=0

Answer by stanbon(75887)

what is the slope of each line and then determine whether the two lines are parallel, perpendicular or neither
-5x+3y=6
y = (5/3)x + 2
slope = 5/3
------------------
-20x+12y=0
y = (20/12)x
y = (5/3)x
slope = 5/3
------------------
Lines are parallel
Cheers,
Stan H.